### v6

# Testing the goodness-of-fit of a functional autoregressive model W. González–Manteiga1, M.D. Ruiz–Medina2, M. Febrero–Bande1

Abstract: The proposed Goodness–of–Fit (GoF) test for checking the linear autocorrelation model in a functional time series is based on an empirical process, whose residual marks and covariate index set are in a separable Hilbert space A functional central limit theorem is derived providing the convergence of the empirical process to a time-changed Wiener process evaluated in a separable Hilbert space with subordinator given by the marginal probability of the involved Autoregressive Hilbertian process (AR(1) process). The large sample behavior of the test statistics is obtained under simple and composite null hypotheses. The consistency of the test is addressed under simple null hypothesis. The simulation study provided in the Appendix illustrates the finite–sample performance of the testing procedure under different families of alternatives.
Key words and phrases: AR(1) process; generalized functional empirical processes; goodness–of–fit test.
1 Introduction
Functional time series models have been extensively analyzed in the last few decades, supporting inference on stochastic processes. Special attention has been paid to functional time series defined in terms of random functions observed through time in regular periods. One can refer the reader to the monograph by Bosq (2000) for linear functional time series in a state space framework. There are several motivating examples arising in Demography, Finance, Environmental Sciences and many other contexts. Weakly dependent functional times series analysis (see, e.g., Hörmann and Kokoszka (2010)) constitutes a first methodological attempt in the literature for the extension of the classical vector–valued time series framework. Also it provides a more flexible analysis than the one based on independent and identically distributed functional random variables (see Horváth and Kokoszka (2012) on functional data analysis, and Ferraty and Vieu (2006) on nonparametric functional statistics).
Long run covariance estimation has played a crucial role in functional time series (see Berkes et al. (2016)), particularly, in recent developments in long range dependence analysis (see, e.g., Li et al. (2020); Ruiz-Medina (2022)). In a more general context, one can mention linear functional regression from correlated in time bivariate functional (surface) data (see, e.g., Ovalle-Muñoz and Ruiz-Medina (2024); Ruiz-Medina et al. (2019)).
In the last few years, there exists a vast literature on testing different hypothesis on the dependence structure in functional time series. In particular, testing independence constitutes a key subject in this research area (see, e.g., Hlávka et al. (2021); Horváth et al. (2013); Zhang (2016)). On the other hand, second–order properties like separability, stationarity and linear correlation range are usually tested with the objective of the specification of the covariance operator family adapted for these models (see Constantinou et al. (2018); Horváth et al. (2014); Kokoszka and Reimherr (2013); Kokoszka et al. (2017); Zhang and Shao (2015), among others). Alternative contributions can be found in testing the assumption of normality and periodicity, as given in Górecki et al. (2018) and Hörmann et al. (2018). The issue of change point testing of the mean, of the variance or of the autocorrelation operator also leads to important contributions, as it can be seen in the already mentioned book by Horváth and Kokoszka (2012). In all the described procedures, the calibration of the distributions of the associated statistical tests has been achieved from the asymptotic theory or resampling techniques, as for example the bootstrap.
A very recent topic, of high interest, in the context of functional time series, is the goodness of fit (GoF) of the structure of the models to be used in one omnibus way. A clear motivation is given by the recent availability of high frequency data in several fields, as for example in finance, where a simple structure for the model can be useful for the forecasting. The recent papers by Amato et al. (2024) and Elías et al. (2022) constitute some examples where statistical analysis based on FAR (Functional Autoregressive) and AR (Autoregressive Hilbertian) models is considered, providing a generalization of the well–known AR (autoregressive) time series setting widely used in forecasting.
The GoF setup for regression models in the Euclidean context is a well–established field with several contributions in the last decades. For a comprehensive review, see González-Manteiga and Crujeiras (2013). A good review for the GoF of regression models with functional data, under the scenario of independent and identically distributed functional random variables, can be found in González-Manteiga (2022). However, there are few papers devoted to specification tests for functional time series. They mainly cover with noise testing, i.e., the case of functional white noise null hypothesis (see, for example, Zhang (2016), Kim et al. (2023) or more recently Kim et al. (2024)), with generalizations to the GoF for functional autoregressive models (FAR), using different approximations based on the periodogram operator in the spectral domain, or the autocovariance estimates constructed from principal components projections.
The present paper extends in the linear setting the methodological GoF approach in Koul and Stute (1999), based on empirical processes, to the context of functional autoregressive models. Specifically, our test statistics involves an empirical process marked by functional residuals, and indexed by a Hilbert–valued covariate. This approach has also been adopted in Cuesta-Albertos et al. (2019) in the context of the functional linear model with scalar response. See also the computational approaches implemented in Álvarez Liébana et al. (2025) and García-Portugués et al. (2021) for the case of funtional response and covariate.
A challenging topic in the above referred GoF functional testing framework is the derivation of the limiting process characterizing the asymptotic properties of the involved generalized –valued empirical process. Theorem 2 of the present paper solves this open problem. Specifically, applying an invariance principle based on Robbins-Monro procedure (see Theorem 2 in Walk (1977)), we prove the convergence in distribution, under simple and composite null hypothesis of our test statistics to an –valued Wiener process, whose index set is also The consistency of the test is proved under simple null hypothesis in Section 3.3.
The large sample analysis of the test, performed under composite is based on the strong–consistency results derived in Chapter 8 in Bosq (2000) when the eigenvalues of the autocovariance operator of an AR(1) process are unknown, covering the cases of known and unknown eigenfunctions. These results allow us to prove the asymptotic equivalence in probability of the test statistics under simple and composite (see Theorem 3).
The asymptotic power analysis deserves further attention and will be addressed in a subsequent paper. The finite–sample power properties are illustrated in the simulation study undertaken, under simple and composite null hypothesis, considering two different alternative hypothesis scenarios. Specifically, Section 2.1 of the Appendix considers nonlinearities in the functional covariates defining the alternative, assuming the eigenvalues and eigenfunctions of the autocovariance operator are unknown under composite While Section 2.2 in the Appendix illustrates the performance of the GoF testing procedure to discriminate between two families of linear autocorrelation operators, in the framework of SPAR(1) models. Here, under composite assuming invariance of the involved operators, the eigenfunctions are known and the eigenvalues are unknown. Under both scenarios, for relatively small functional sample sizes, robust empirical test sizes, and good empirical power properties are observed under simple The impact of the misspecification level of the autocorrelation operator is also illustrated under composite As expected, when the misspecification level is lower, that is the case of invariant SPAR(1) model scenario, where the eigenfunctions are known, but the eigenvalues are unknown, better performance is observed for small sample sizes than in the case of unknown eigenfunctions and eigenvalues analyzed in Section 2.1 of the Appendix.
The outline of the paper is the following. Preliminaries are provided in Section 2. The formulation of the GoF testing procedure and its implementation in practice are given in Section 2.1. The asymptotic distribution of the marginals of the generalized –valued empirical process is provided in the Central Limit Theorem derived in Section 3.1.The Functional Central Limit Theorem in Section 3.2 introduces the limiting generalized –valued Gaussian process, which is identified, in probability distribution, with time–changed –valued Wiener process under From Continuous Mapping Theorem, the critical value for the supremum norm of the empirical standardized random projected test statistics is obtained from the standard Gaussian distribution, by applying reflection principle given in terms of the boundary crossing probabilities of Brownian motion. Consistency of the GoF test is addressed in Section 3.3. The composite scenario is analyzed in Section 4. Final discussion is presented in Section 5. Auxiliary information is provided in the Appendix. Also, illustration of the finite sample performance of the proposed testing procedure is provided in the simulation study undertaken in this Appendix.
2 Preliminaries
In what follows we denote by the space of zero–mean –valued random variables on the basic probability space with Let be a zero–mean AR(1) process on the basic probability space satisfying
where denotes the autocorrelation operator of AR(1) process with being the space of linear bounded operators on Along the paper, we assume that the conditions formulated in Chapter 3 in Bosq (2000) hold, ensuring the existence of a unique stationary solution to equation (2.1). Hence, denotes the autocovariance operator of and its cross–covariance operator.
The –valued innovation process is assumed to be –valued Strong White Noise (–SWN) (see Definition 3.1 in Bosq (2000)). That is, is a sequence of independent and identically distributed zero–mean –valued random variables with for all and functional variance that coincides with the trace norm of the autocovariance operator of We restrict our attention to AR(1) processes satisfying the following condition:
Assumption A1. is a strictly stationary AR(1) process with –SWN innovation such that
Under model (2.1), keeping in mind Assumption A1, the following orthogonality condition
holds for with denoting the almost surely (a.s.) equality.
In what follows, we will use the notation and for its indicator function, and
Let us consider a centered non–degenerated Gaussian probability measure on From Theorem 4.1 in Cuesta-Albertos et al. (2007), under Assumption A1, if satisfies Carleman condition, a class in of Borel sets of defines a separating class of if it contains a set of positive –measure. Thus, Lemma 1(d) in Escanciano (2006) can be reformulated in an infinite–dimensional framework by considering, for each and for a given orthonormal basis of
leading to the definition of the separating class of sets (see Theorem 1.2.1 in Prato and Zabczyk (2002)). Hence, under Assumption A1, one can equivalently express the orthogonality condition (2.2) as
for a given orthonormal basis of It is straightforward that the sequence is an –valued martingale difference with respect to the filtration where for as given in the following lemma:
Lemma 1.
For each the sequence
is an –valued martingale difference with respect to the filtration where for
The proof is straightforward from for every with being the orthogonal projector into (see equation (2.55), and pp. 72–73 in Bosq (2000)).
Along the paper, from (2.1), the following identity will be applied several times (see equation (3.11) in Bosq (2000)), for every
2.1 Testing the autocorrelation model
We first address the functional parametric testing problem
where is a known operator. A Kolmogorov–Smirnov (K–S) test based on the following Hilbert–valued marked empirical process indexed by is adopted
where is an orthonormal basis of the separable Hilbert space and denotes the autoregressive process introduced in equation (2.1). As before, for each
From Theorem 2 below (see Section 3.2), weak converges to an –valued Gaussian process indexed by with covariance operator for every where for and
Thus, Here, denotes the infinite–dimensional vector
Then, can be identified in distribution sense with time–changed –valued Wiener process with subordinator (see Remark 1 below).
In practice, applying random projection methodology, assuming satisfies Carleman condition (see Theorem 4.1 in Cuesta-Albertos et al. (2007)), the null hypothesis will be rejected if conditionally to the observation of an value of a functional non–degenerated Gaussian random variable, the test statistic
exceeds a critical value of the probability distribution of the supremum norm of Brownian motion on the interval which is computed by Reflection Principle from the standard normal probability distribution as In equation (2.8),
and denotes the infinite–dimensional marginal probability measure induced by with
An alternative subordinator scenario is discussed in Section 1.3 of the Appendix.
In the computation of the second–order moments of –valued martingale difference sequence (see Section 1.2 of the Appendix), the following assumption will be considered:
Assumption A2. Assume that is independent of for all
3 Asymptotic properties for simple hypothesis ()
The asymptotic distribution of the generalized –valued empirical process in equation (2.5) when (simple hypothesis) is analyzed in this section. Specifically, Theorem 1 in Section 3.1 provides the convergence in distribution as where, for each Here, denotes the zero–mean Gaussian distribution on with autocovariance operator Theorem 2 in Section 3.2 then obtains the weak convergence of to a generalized –valued Gaussian process identified in distribution sense with time–changed –valued Wiener process with subordinator having covariance operator
3.1 Central Limit Theorem
Theorem 2.16 in Bosq (2000) is now applied in the derivation of Theorem 1, providing the Gaussian limit distribution of the marginals of the generalized –valued marked empirical process in (2.5).
Theorem 1.
Let be a centered AR(1) process. Under Assumptions A1–A2, for each the generalized –valued empirical process (2.5) satisfies as with denoting the convergence in distribution. Here, for each the autocovariance operator
If, for every
Proof.
The proof is based on the verification of the conditions assumed in the CLT for Hilbert–valued martingale difference sequences given in Theorem 2.16 in Bosq (2000). Let us first verify condition (2.59) in Theorem 2.16. For every and the following events are considered:
Clearly, and for every
From equation (3.12), applying Chebyshev inequality and stationarity, keeping in mind that the events and coincide for and we obtain
Dominated Convergence Theorem yields to
On the other hand, for any and applying SWN property of the innovation process under Assumptions A1–A2, from equation (2.4),
Thus, is uniformly integrable, and the desired result holds.
Let us now prove that condition (2.61) in Theorem 2.16 in Bosq (2000) holds. Specifically, considering again Assumptions A1–A2, from equation (2.4), and from the SWN property of applying Markov Theorem, for any
From (3.16), in view of the trace property of operator
as we wanted to prove.
Finally, we prove condition (2.60) in Theorem 2.16 in Bosq (2000) holds. Specifically, we prove that condition (2.36) in Corollary 2.3 in Bosq (2000) holds. Under Assumptions A1–A2, considering, as before,
with
we have
Let us now denote for with, as before, for every
Under Assumptions A1–A2, applying strictly stationarity, and that is strong white noise
Note that under Assumption A1, and, since is a trace operator, its Hilbert–Schmidt operator norm is finite. Thus, we can apply Corollary 2.3 in Bosq (2000) with and, for all and for each
leading to condition (2.60) in Theorem 2.16 in Bosq (2000), as we wanted to prove. Thus, the weak convergence to the indicated –valued Gaussian random variable holds.
3.2 Functional Central Limit Theorem
This section provides the convergence in law of the generalized (indexed by ) functional empirical process (2.5) to a generalized -valued Gaussian process. Continuous Mapping Theorem, and Theorem 4.1 in Cuesta-Albertos et al. (2007) will then be considered to formulate the distribution free test (2.8) (see Section 2.1).
The following lemma (see Theorem 2 in Walk (1977)), providing an invariance principle based on Robbins-Monro procedure, will be applied in the proof of Theorem 2 below.
Lemma 2.
Let be a martingale difference sequence of –valued random variables, with respect to the filtration satisfying Let be a trace operator. For each denote as the autocovariance operator of given That is,
Assume that
For
Then, the sequence of random elements in with the supremum norm, which are defined by
converges in distribution to a Brownian motion in with a.s., and covariance operator of
The following additional condition is required in the derivation of the next result.
Assumption A3. where denotes the sample functional value in corresponding to of at time
In what follows will denote the separable Banach space of –valued continuous functions on with respect to the norm, under the supremum norm
Theorem 2.
Remark 1.
The limiting process can be identified in probability distribution sense with time–changed –valued Brownian motion with subordinator since, from equation (3.21), for every Here, process is an infinite–dimensional (–valued) Wiener process on the interval in the sense introduced in Definition 2 of Dedecker and Merlevéde (2003). In particular, induces a probability measure on The identity has been considered in the determination of the critical value of the test statistics (2.8) when we apply random projection methodology. An alternative subordinator scheme has been considered in Section 1.3 of the Appendix.
Proof.
Let us define the sequence satisfying
for each where is the –valued empirical process introduced in equation (2.5).
The proof is based on verifying that, for each fixed the –valued martingale difference sequence satisfies Lemma 2(i)–(iii), after proving that, for every
Note that, applying Chebyshev’s inequality, for
where , and We now verify conditions (i)-(iii) of Lemma 2, ensuring ergodicity of the involved centered stationary martingale difference functional sequence. We begin considering (i) in Lemma 2. Thus, under Assumption A1, keeping in mind (2.4), for and
Lemma 2(ii) holds, under Assumptions A1–A2, from equation (1.37), since, for each and every
Finally, Lemma 2(iii) also holds under Assumptions A1–A3. Specifically, we prove convergence in Thus, from equation (2.4), for each and applying Chebyshev’s inequality, we obtain as
under Assumption A1 and A3, since ∎
3.3 Consistency
This section shows the consistency under Assumption A1 of the proposed GoF procedure based on for testing versus the alternative when is a known operator. By it should be understood that Let for Assume that for every if and only if Let for every and consider For and let also where and means expectation is computed under the alternative (). Then, under Assumption A1, applying stationarity, for every
in view of the short range dependence displayed by AR(1) process From Corollary 2.3 in Bosq (2000), equation (3.25) leads to the almost surely convergence
yielding the consistency of the test.
4 Asymptotic properties for composite hypothesis
This section addresses the asymptotic analysis under composite null hypothesis. Specifically, we consider that the probabilistic conditions, determining the scenario of the null hypothesis, ensure consistency of an estimator of the unknown Two scenarios are considered respectively corresponding to the cases where the eigenfunctions of the autocovariance operator of are known and unknown.
Let us first consider the composite null hypothesis
where the family of autocorrelation operators in satisfies the conditions in Lemma 8.1, and Theorems 8.5–8.6 in Bosq (2000). Thus, we derive the asymptotic properties of the generalized –valued plug–in empirical process
where is a consistent estimator of under in the norm of the space of bounded linear operators on (see Bosq (2000)).
In this first scenario, when the eigenfunctions of the autocovariance operator of are known, the estimator of is given by
where and and
The following assumption is also considered providing the strong consistency of in (see Bosq (2000)).
Assumption A4. Assume the following conditions:
is a standard AR(1) process with –SWN innovations, and
The eigenvalues of satisfying for every are strictly positive (i.e., ), and the eigenfunctions are known.
for every
Under Assumption A4, is bounded satisfying: with
Considering, additionally, the conditions assumed in Lemma 8.1, and Theorems 8.5–8.6 in Bosq (2000), we obtain the the strong consistency of in the space
Let us now consider the composite null hypothesis
where the family of autocorrelation operators in satisfies the conditions in Theorems 8.7–8.8 in Bosq (2000). The estimator of is formulated in terms of the empirical eigenfunctions and eigenvalues given by
where Specifically, is defined as
with, as before, and and for and
The following additional condition is assumed for strong–consistency (see Theorems 8.7–8.8 in Bosq (2000)).
Assumption A5. The AR(1) process is such that
The eigenvalues of satisfy
where as before, for every
For every and a.s.
Under the assumed conditions (see Assumption A5), Theorems 8.7–8.8 in Bosq (2000) establishe the strong consistency of in in the case of unknown eigenfunctions of the autocovariance operator Thiese theorems are applied in the derivation of the next result providing the asymptotic equivalence in probability of and
Theorem 3.
Under conditions of Theorems 8.7–8.8 in Bosq (2000), the following identity holds:
Proof.
Let us consider, for every
From Theorems 8.7–8.8 in Bosq (2000), Also, from equation (2.21) in Theorem 2.5(2) in Bosq (2000), in a similar way to the proof of Theorem 3.9 (see, in particular, equations (3.36)–(3.37)) in Bosq (2000), Thus, (4.32) implies equation (4.31).
Remark 2.
The asymptotic equivalence in probability of and also follows from the strong–consistency in of
As direct consequence of Theorem 3, under Assumptions A1–A5, and conditions of Theorems 8.7–8.8 in Bosq (2000), as and From Remark 2, similar assertions hold regarding the limiting generalized –valued Gaussian process of the empirical process The remaining steps in the implementation of the testing procedure via random projection follows in a similar way to Section 2.1, where can be defined as in equation (2.9) in terms of the plug–in empirical process (respectively, the plug–in empirical process under the first scenario).
5 Discussion
Up to our knowledge, this paper constitutes the first attempt to derive asymptotic theory of GoF test of the AR(1) time series model, based on the empirical process methodology. The main contribution of this paper is the functional central limit result derived. The analysis under composite null hypothesis is supported by the strong–consistency results in Chapter 8 of Bosq (2000). Note that the asymptotic analysis performed here is different from the asymptotic GoF analysis of the classical functional linear model under independent functional data (see Section 1.1 of the Appendix).
The performance of our approach is also illustrated in the context of SPAR(1) processes in Section 2.2 (see, e.g., Caponera and Marinucci (2021)). The assumed invariance property of the kernels defining the autoregression, and covariance operators of SPAR(1) process leads to an important dimension reduction in the implementation. Similar results hold under this invariance property in the case of autoregressive processes on compact and connected two point homogeneous spaces (see, e.g., Ma and Malyarenko (2020)).
Acknowledgements
This work has been supported in part by projects PID2022–142900NB-I00 and PID2020-116587GB-I00, financed by MCIU/AEI/10.13039/501100011033 and by FEDER UE, and CEX2020-001105-M MCIN/AEI/10.13039/501100011033), as well as supported by grant ED431C 2021/24 (Grupos competitivos) financed by Xunta de Galicia through European Regional Development Funds (ERDF).
Appendix
Auxiliary information about differences arising in our setting with respect to the asymptotic theory of GoF in the –valued linear model, under independent functional data, are discussed. The second–order properties of –valued martingale difference sequence are provided. An alternative subordination scheme, in the spirit of the real–valued case, is introduced. The finite sample performance of the GoF testing procedure proposed is finally illustrated in the simulation study undertaken.
1 Auxiliary information
Additional supporting information to facilitate the reading of the paper is provided in this section.
1.1 Differences with asymptotic GoF under independent data
Theorem 1 in Cardot et al. (2007) focuses on functional regression with scalar response, and –valued covariate The projection estimator of the regression function does not satisfy a Central Limit Theorem (CLT) with convergence to a non–degenerated random element in Weak–convergence results hold for the predictor, under suitable truncation and convexity of the eigenvalues of the autocovariance operator of the covariate In the AR(1) framework, the situation is quite different as follows from the convergence results enumerated below (see Mas (1999)):
The convergence to zero, in probability, of in the norm of the space of Hilbert–Schmidt operators on , under suitable conditions like the ones assumed in Theorem 4.1, p. 98, in Bosq (2000). Here, denotes the projection operator into the subspace of generated by the eigenfunctions of In particular, under the conditions assumed in Theorem 4.1 in Bosq (2000), our choice of the truncation order must be such that, as ensuring, in particular, Proposition 4 in Mas (1999) holds. Note that Theorem 4.1 in Bosq (2000) is proved by applying the strong law of large number for weakly dependent sequences of -valued random variables, that leads to the strong law of large numbers for AR(1) processes (see, e.g., Theorem 3.7, p.86, in Bosq (2000)). Specifically, from Lemma 4.1, p.96, in Bosq (2000), on the AR representation of the diagonal self–tensorial product of an AR(1) process, the strong consistency in the norm of of the empirical autocovariance operator of is obtained by applying Theorem 3.7 in Bosq (2000).
The convergence in distribution of with to a centered Gaussian random Hilbert-Schmidt operator under the key condition This limit result is obtained by applying a CLT for an array of –valued martingale differences. The limit centered Gaussian random Hilbert-Schmidt operator has covariance operator defined by, for
(1.35) Note that, in the functional linear regression model with scalar response and functional covariate, the AR(1) condition can not be considered since the innovation process is real–valued. Hence, the convergence to a non–degenerated random element in the norm of does not hold (see Theorem 1 in Cardot et al. (2007)). A similar assertion can be made for the functional linear regression model with functional response and covariate (see Theorem 8 in Crambes and Mas (2013)), since the above–referred CLT for an array of –valued martingale differences can not be applied under independent functional data.
Condition means that Thus, or, equivalently, belongs to the Reproducing Kernel Hilbert Space (RKHS) generated by the integral operator In the case where and have a common system of eigenfunctions, this condition can be equivalently expressed as
Hence, a faster decay of the eigenvalues of the autocovariance operator of the innovation process than the eigenvalues of the square autocovariance operator of the AR(1) process is required. In particular, a mild sufficient condition is given by
Section 2.3 in Mas (1999) provides some rules to compute the truncation order depending on the functional sample size in order to ensure the derived asymptotic results hold. Specifically, one can consider if obeys the following asymptotic behavior as On the other hand, for one can choose if while if
In the case where the eigenfunctions of the autocovariance operator are unknown, projection into the empirical eigenfunctions of can be considered, and a CLT is obtained in Theorem 8.9 in Bosq (2000). Indeed, a similar decomposition to equation (11) in Cardot et al. (2007) can be obtained. The strong consistency of the empirical eigenvalues and eigenfunctions of is then applied (see Theorem 4.4, Lemma 4.3, Theorem 4.5, and Corollary 4.3 in Bosq (2000)).
The decay velocity of the eigenvalues of and the distance between such eigenvalues, characterized by their distribution, mainly in the interval play a key role in the conditions assumed for the uniform asymptotic equivalence in probability in the norm of of the totally specified (under simple ), and plug–in empirical processes (under the two composite null hypotheses considered).
1.2 Second–order properties of involved –valued martingale difference sequence
Note that, for each the strongly integrability of the marginals of the –valued martingale difference sequence allows the computation of the means and conditional means of the elements of this sequence from their weak counterparts (Section 1.3 in Bosq (2000)).
Under Assumptions A1–A2, for any and and can be computed as follows:
where we have applied for all under Assumptions A1–A2. As before, denotes the infinite–dimensional marginal probability measure induced by and is the event for each Keeping in mind equation (2.4), applying the strictly stationarity of for each the autocovariance operator of is given by
In a similar way, the covariance operator can be computed for every
where is interpreted as in equation (2.7) of the paper. Here, for if and if
1.3 Subordinators in the probability distribution identification of
This section provides an alternative subordinator scenario to work in the design of the test statistics based on random projection methodology (see Section 2.1 of the paper). Specifically, when indentifiability is performed in the norm of the space of nuclear operators the spectral properties of the autocovariance operator of the innovation process are summarized in terms of its trace norm, providing the functional variance of the residual marks of the empirical process in equation (2.5) of the paper, in the spirit of the real–valued case (see Koul and Stute (1999)). One can then consider for every
with, as in the paper,
Note that is a nondecreasing and nonnegative function, whose values are in the interval with One can then consider the identification in probability distribution of the limiting process with time–changed –valued Wiener process with subordinator In particular, is independent of and defines a Gaussian distribution on with zero mean and covariance operator
Here,
2 Illustration of the performance of GoF
This section presents two illustrations of the proposed GoF test in the context of AR(1) (), and SPAR(1) models in Sections 2.1 and 2.2, respectively.
2.1 Detecting independence and nonlinearities
In this section, an illustration of the performance of the proposed GoF procedure is given by simulation in both scenarios, simple and composite hypothesis. For simple hypothesis, we consider the particular case of testing versus in the AR(1) model (2.1). Under simple the empirical type I error and power, based on repetitions of the procedure implemented, for functional samples sizes are computed. The critical values of the test statistics are approximated from Fast Bootstrap based on bootstrap replicates. As given in Section 2.1 (see also Section 3.2), the random projection methodology is implemented in such computations (see Cuesta-Albertos et al. (2019)).
Under a Gaussian scenario, the functional values of an AR(1) process are generated with support in the interval evaluated at temporal nodes. A Gaussian zero–mean –SWN noise innovation process having integral autocovariance operator with exponential covariance kernel
is generated. A Gaussian random initial condition is also generated, independently of the innovation process . has exponential autocovariance kernel having the same scale parameter values, as the innovation process . Equation (2.1) is then recursively generated involving the numerical approximation of the integral where, under the alternative hypothesis , the kernel of the integral autocorrelation operator is given by for . The constants are ensuring that the process is stationary. To remove dependence from the random initial condition, a burn-in period of observations is considered. That is, for each functional sample size the above–described recursive approximation of the values is computed removing the values
The random projection of the functional covariate and marks is obtained. The empirical process in equation (2.5) is then expressed in terms of such random projections as
where and are independent realizations of a zero–mean Gaussian process with covariance kernel given by the autocovariance kernel of and being those processes, and approximated by their truncated, at term Karhunen–Loéve expansions. Note that here we have considered two independent realizations of an –valued Gaussian random variable for random projection, but, in our particular AR(1) scenario, it is suffcient to consider for random projection a common realization of a non–degenerated –valued Gaussian random variable. Fast Bootstrap is implemented from the equation
where are independent and identically distributed standard normal random variables. The computation of the –value is obtained for each projection from
where, as commented, bootstrap replicates have been considered. The numbers of projections tested is To obtain only one –value the False Discovery Rate is computed (i.e., the expected proportion of false positives among the rejected hypotheses).
For the case of composite null hypothesis the process generated is an AR(1) process, that uses the same employed to the alternative hypothesis in the case of simple null hypothesis. For the alternative, we have also considered an AR(1) process with the same but evaluated in the square process, i.e.
where the parameter avoids that the process escapes to nonstationarity, being only when . Although, this is a clearly non AR(1) process, it is pretty close to an ARH(1) due to the scale of . Now, the statistics in equation (2.41) must change to incorporate the variability due to the estimation of parameter . So, following Escanciano (2006), is changed by given by
where is the estimation of the functional parameter and with being the resampled errors from the estimated model.
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.050 | 0.054 | 0.044 | 0.038 | 0.034 | 0.034 | 0.042 | |
| 0.060 | 0.056 | 0.054 | 0.062 | 0.058 | 0.060 | 0.056 | |
| 0.050 | 0.054 | 0.056 | 0.052 | 0.048 | 0.052 | 0.040 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.434 | 0.492 | 0.526 | 0.524 | 0.530 | 0.562 | 0.552 | |
| 0.486 | 0.630 | 0.678 | 0.728 | 0.764 | 0.836 | 0.848 | |
| 0.650 | 0.808 | 0.896 | 0.922 | 0.932 | 0.984 | 0.994 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.054 | 0.034 | 0.034 | 0.048 | 0.056 | 0.038 | 0.032 | |
| 0.056 | 0.062 | 0.058 | 0.050 | 0.068 | 0.046 | 0.046 | |
| 0.050 | 0.050 | 0.052 | 0.046 | 0.054 | 0.048 | 0.048 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.114 | 0.096 | 0.084 | 0.062 | 0.064 | 0.062 | 0.066 | |
| 0.152 | 0.182 | 0.196 | 0.204 | 0.214 | 0.232 | 0.264 | |
| 0.242 | 0.340 | 0.378 | 0.400 | 0.432 | 0.520 | 0.508 | |
| 0.302 | 0.438 | 0.528 | 0.562 | 0.584 | 0.666 | 0.706 | |
| 0.448 | 0.610 | 0.680 | 0.756 | 0.806 | 0.892 | 0.938 | |
| 0.474 | 0.682 | 0.774 | 0.830 | 0.888 | 0.970 | 0.992 |
In Tables 1 and 2, one can respectively find the –values and power approximations computed, based on repetitions of the implemented testing procedure, for functional sample sizes The Confidence Interval, based on repetitions, obtained for is Specifically, Table 1 shows the rejection probability when is true. One can observe in Table 1 that the computed estimates of the test size are slightly below the true parameter value for the functional sample size . While, for functional sample sizes an improvement is observed in the computed estimates of for all number of random projections. For the functional sample sizes analyzed, no patterns are observed regarding computed estimates and the number of random projections. The empirical power of the test, for the same functional sample sizes and number of repetitions, are displayed in Table 2. As expected, the performance of the test is improved when the functional sample size increases. Tables 3 and 4 show sizes and powers for the composite hypothesis. The test seems well calibrated as every element in Table 3 is inside the Confidence Interval for although, for the test shows a tendency to be below similar to what happens in simple null hypothesis. Given the rate of convergence of the empirical eigenfunctions to the theoretical ones (see Lemma 4.3 and Theorem 4.5 in Bosq (2000)), larger functional sample sizes are required to obtain competitive empirical power values. In Table 4, we illustrate this fact considering additional functional sample size values
2.2 Spherical functional time series SPAR(1)
In this section we analyze the performance of the proposed GoF procedure in the special case where is the space of real–valued square integrable functions on the –dimensional sphere in Here, denotes the normalized Riemannian measure on In what follows, we will denote by the orthonormal basis of eigenfunctions of the Laplace–Beltrami operator on with denoting the dimension of the th eigenspace of the Laplace–Beltrami operator. Here, we consider and
The time–varying random projections of the SPAR(1) process, with repect to the eigenfunctions of the Laplace Beltrami operator, have been generated by using the MatLab function arima. They are evaluated at time instants, where denotes the functional sample size. The functional values of the innovation process are also generated from its time–varying projections with respect to such a basis at times.
In the generations of SPAR(1) model we have considered the spherical harmonics which are localized in the first four eigenspaces () of the Laplace Beltrami operator on (see left–hand–side of Figure 1). Under the simple null hypothesis one realization (evaluated at spherical spatial nodes) of the SPAR(1) process projected into the corresponding direct sum of Laplace Beltrami eigenspaces is displayed at the right–hand–side of Figure 1. Here, denotes the th eigenspace of the Laplace Beltrami operator on The eigenvalues of the invariant kernel defining the autoregressive operator under simple associated with the Laplace Beltrami eigenfunctions, are given by
We first consider versus where we assume that the response is of the form with for and being an –SWN. Thus, the random components of the innovation process still display significative correlations induced by Both operators, and are invariant bounded linear operators against the group of rotations in the sphere. The test statistics (2.8) is evaluated at different random projections generated from Fractional Brownian Motion (FBM). In the generation of FBM we have used MatLab function wfbm. The critical value corresponding to is for the supremum norm of Brownian motion on the interval As indicated, this value is computed, by applying Reflection Principle, from the standard normal probability distribution (see Section 2.1).
Under simple null hypothesis (totally specified autocorrelation operator), Table 5 displays the empirical test size for the functional sample sizes Again, no patterns are displayed. In particular, no tendency to be below is observed for In Table 6, one can find the empirical test power values, based on repetitions, for the functional sample sizes As in the previous scenario analyzed in Table 2, one can observe in Table 6 a better performance of the test statistics when the functional sample size increases.
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.060 | 0.040 | 0.040 | 0.060 | 0.040 | 0.040 | 0.060 | |
| 0.040 | 0.040 | 0.040 | 0.060 | 0.060 | 0.060 | 0.040 | |
| 0.050 | 0.054 | 0.056 | 0.052 | 0.048 | 0.052 | 0.040 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.376 | 0.336 | 0.718 | 0.454 | 0.542 | 0.594 | 0.584 | |
| 0.566 | 0.464 | 0.658 | 0.700 | 0.428 | 0.450 | 0.762 | |
| 0.562 | 0.772 | 0.806 | 0.772 | 0.744 | 0.882 | 0.732 | |
| 0.838 | 0.546 | 0.682 | 0.944 | 0.864 | 0.912 | 0.892 | |
| 0.890 | 0.998 | 0.914 | 0.994 | 0.978 | 0.984 | 0.984 | |
| 0.851 | 0.926 | 0.974 | 0.938 | 0.850 | 0.996 | 0.992 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.018 | 0.014 | 0.014 | 0.022 | 0.028 | 0.022 | 0.022 | |
| 0.032 | 0.034 | 0.026 | 0.022 | 0.034 | 0.030 | 0.024 | |
| 0.038 | 0.052 | 0.032 | 0.042 | 0.040 | 0.038 | 0.040 | |
| 0.042 | 0.054 | 0.032 | 0.026 | 0.056 | 0.040 | 0.042 | |
| 0.046 | 0.046 | 0.050 | 0.054 | 0.046 | 0.044 | 0.044 | |
| 0.050 | 0.045 | 0.050 | 0.045 | 0.045 | 0.045 | 0.055 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.636 | 0.706 | 0.522 | 0.506 | 0.484 | 0.584 | 0.576 | |
| 0.668 | 0.736 | 0.630 | 0.652 | 0.556 | 0.806 | 0.678 | |
| 0.824 | 0.870 | 0.646 | 0.912 | 0.612 | 0.062 | 0.638 | |
| 0.962 | 0.826 | 0.670 | 0.902 | 0.882 | 0.914 | 0.764 | |
| 0.994 | 0.966 | 0.998 | 0.978 | 0.986 | 0.918 | 0.840 | |
| 0.998 | 0.986 | 0.996 | 0.826 | 0.786 | 0.976 | 0.960 |
Let us now consider, under the general scenario given by condition (c0) in Bosq (2000), ensuring existence and uniqueness of a stationary solution to the AR(1) equation (2.1) (see Bosq (2000), p.74, Chapter 3), the following testing problem:
where generations under have been achieved for with eigenvalues Hence, under Generations under have been performed, as before, under the pure point spectrum ensuring is satisfied, since Note that Lemma 8.1(3) in Bosq (2000) holds under the generated scenario, since for and computed from terms of the empirical sequence Here, denotes the nearest positive integer.
Tables 7 and 8 show empirical test sizes and powers for composite null hypothesis under a SPAR(1) scenario. For the number of random projections, functional sample sizes and number of repetitions reflected in Tables 7 and 8, the values of the test statistics (2.8) based on the plug–in empirical process (4.27) are respectively computed under and under The eigenvalues of are estimated from the following identity (see equation (3.13) in Bosq (2000)):
where denotes the adjoint of with in the generations. Given the diagonal spectral factorization in terms of the common resolution of the identity of operators in equation (2.43), the associated pure point spectra satisfy
Equation (2.44) is then approximated in terms of the following empirical pure point spectra:
where, for and denote the generated repetitions of and under respectively, for the estimation of under in the SPAR(1) model. Note that, from equations (2.44)–(2.45), considering in the a.s. convergence in –norm in Lemma 8.1(3) in Bosq (2000), the a.s. convergence of the eigenvalues of (computed from (2.44)–(2.45)) to the eigenvalues of satisfying (2.44) is obtained.
In Table 7, we observe empirical test sizes close to the theoretical value for the sample sizes The empirical test powers showed in Table 8 are also computed for the functional sample sizes considering the same number of repetitions As expected, the empirical test powers are improved when the functional sample size increases. A better performance in the case of composite hypothesis is observed under a lower level of misspecification of the autocorrelation operator for small sample sizes (). That is the case of the SPAR(1) scenario, where the eigenvalues are unknown but the eigenfunctions are known, in contrast with the scenario where the eigenfunctions and eigenvalues are unknown analyzed in Table 4 for Under simple the reverse situation is observed in Tables 2 and 6. Although, given the variability displayed by columns in these two tables, for the sample sizes one can not conclude that these numerical results establish any significative performance difference between the two testing problems.