### v7

# Testing the goodness-of-fit of a functional autoregressive model

Abstract
The proposed Goodness–of–Fit (GoF) test for checking the linear autocorrelation model in a functional time series is based on an empirical process, whose residual marks and covariate index set are in a separable Hilbert space A functional central limit theorem is derived providing the convergence of the empirical process to a time–changed Wiener process evaluated in a separable Hilbert space with subordinator given by the marginal probability of the involved strictly stationary Autoregressive Hilbertian process (AR(1) process). The large sample behavior of the test statistics is obtained under simple and composite null hypotheses. The consistency of the test is addressed under simple null hypothesis. The finite–sample performance of the testing procedure, under different families of alternatives, and random projection schemes, is illustrated in the Appendix.
keywords:
AR(1) process , generalized functional empirical processes , goodness–of–fit test.organization=1Universidad de Santiago de Compostela
2Universidad de Granada,
Research highlight 1. The infinite-dimensional asymptotic analysis of GoF in the context of AR(1) models applying process subordination
Research highlight 2. The introduction of a new family of testing procedures, based on generalized empirical functional processes, under simple and composite null hypothesis addressing their asymptotic equivalence
1 Introduction
Functional time series models have been extensively analyzed since the’90s, supporting inference on stochastic processes. The monograph by [2] constitutes a benchmarking in asymptotic inference from functional linear time series models in a state space framework. We also refer the reader to [15], [18] on weakly dependent functional times series analysis, and to [10] on nonparametric functional statistics.
Current advances on statistical testing in functional time series cover different issues related to dependence structure, and independence (see, e.g., [13]; [16]; [31]); separability, stationarity and linear correlation range (see [4]; [17]; [21]; [22]; [30]); normality and periodicity (see [12] and [14]). Special attention has been paid to the issue of change point testing of the mean, the variance, and the autocorrelation operator (see [18]).
Specification and Goodness-of-Fit (GoF) tests are emerging areas in the Functional Data Analysis (FDA) literature (see [31], [19], [20], among others). In the context of the functional linear regression model, the asymptotic analysis of GoF tests in infinite–dimensional spaces is currently an open challenge (see, e.g., [5], [1] and [11], and the references therein). Specifically, most of the proposals are based on the application of random projection methodology from Theorem 4.1 in [6], typically using bootstrap techniques for calibration. In the context of functional linear model with scalar response, GoF asymptotic analysis, under simple and composite null hypothesis, is addressed from randomly projected empirical processes in [5]. The GoF asymptotics in the framework of the Functional Linear Model with Functional Response (FLMFR) entails several unsolvable difficulties, also affecting the usual asymptotic analysis based on random projection methodology (see, e.g., [1] and [11]). In this context, Theorem 4.1 in [6] should be formulated in terms of a non–degenerate bivariate infinite–dimensional Gaussian measure. This difficulty is overcoming in the present paper by considering inverse subordination of our test statistics, defined from a Hilbert–valued (–valued) empirical process indexed by an –valued covariate. To this aim, under strictly stationarity, we introduce, in equation (10) below, the analogous to the concept of hitting time, in terms of an infinite–dimensional version of marginal quantile transformation, involving a tight infinite product Gaussian measure (see Theorem 1.2.1 in [7]).
Our proposal extends the GoF asymptotic analysis performed in [23], in the context of real–valued Markovian time series, to the functional time series framework under an Autoregressive Hilbertian process of order one (AR(1) process) scenario. Note that this scenario corresponds to the case of nondecreasing real–valued function being the identity in [23]. In our functional residual marked empirical process based design of the test statistics, we adopt the approach proposed in Lemma 1(d) in [9], reformulating it in an infinite–dimensional framework. Specifically, we apply the fact that any class of Borel sets in containing a set of positive non–degenerate tight infinite–dimensional Gaussian measure defines a separating class (see Theorem 4.1 in [6]). The separating class is then indexed by the functions in the Reproducing Kernel Hilber Space (RKHS) of such an infinite–dimensional Gaussian measure, when the marginals of the AR(1) process satisfy Carleman condition. In particular, the index set of the separating class can be the functions in the RKHS generated by the autocovariance operator of our AR(1) process (see Theorem 1.2.1 in [7]).
The Central Limit Theorem of the present paper provides the asymptotic infinite–dimensional Gaussian probability distribution of the marginals of the generalized functional residual marked empirical process, under simple and composite null hypothesis. Under both null hypothesis scenarios, the limiting process is identified in law with time–changed –valued Brownian motion via the derived Functional Central Limit Theorem (see Theorem 2 in [29]). Continuous Mapping Theorem then allows the asymptotic approximation of the critical values of our infinite–dimensional version of Kolmogorov–Smirnov (K–S) test. In practice, random projection methodology can be applied, and conditionally on a random direction, the weak convergence to real–valued time–changed Brownian motion holds. The critical values are then asymptotically approximated from the boundary crossing probabilities of real–valued Brownian motion on Our application of random projection methodology leads to an asymptotically distribution free (ADF) test statistics. The consistency of the test is proved against a broad class of alternatives, in the infinite–dimensional and conditionally one-dimensional formulations. Note that in the case of misspecified autocorrelation operator, the strong consistency results on the autocorrelation operator estimation obtained in Chapter 8 in [2] play a key role, in our derivation of the asymptotic analysis of our test statistics under composite null hypothesis.
The finite–sample power properties are illustrated in the simulation study undertaken, under simple and composite null hypothesis, considering two different alternative hypothesis scenarios. Specifically, B.1 considers nonlinearities in the functional covariates defining the alternative. The misspecified family of AR(1) models under the null are approximated in a consistent way by projection estimators based on the empirical eigenfunctions and eigenvalues of the autocovariance operator. B.2 addresses the issue in SPAR(1) GoF testing of discrimination between two families of linear autocorrelation operators, under second–order invariance assumption. Thus, the eigenfunctions of the autocovariance operator coincide with the eigenfunctions of the Laplace Beltrami operator in the sphere, leading to the scenario of known eigenfunctions and unknown eigenvalues. Under both scenarios, for relatively small functional sample sizes, robust empirical test sizes, and good empirical power properties are observed under simple The impact of the misspecification level of the autocorrelation operator is also illustrated under composite As expected, when the misspecification level is lower, i.e., invariant SPAR(1) process in B.2, better small sample size performance is observed than in the case of unknown eigenfunctions and eigenvalues considered in B.1.
The outline of the paper is the following. Some preliminary definitions and results are provided in Section 2. In particular, the formulation of the GoF testing procedure, and its implementation in practice are given in Section 2.2. The asymptotic distribution of the marginals, and the limiting process of the generalized functional residual marked empirical process are derived in Section 3. Consistency of the GoF test is addressed in Section 3.3. The misspecified autocorrelation operator scenario is analyzed in Section 4. Final discussion is presented in Section 5. Auxiliary information is provided in the Appendix. Illustration of the finite sample performance of the proposed testing procedure is provided in the simulation study undertaken in this Appendix.
2 Preliminaries
Let be a zero–mean AR(1) process on the basic probability space satisfying the state equation
The autocorrelation operator of is assumed to be in the space of bounded linear operators on The conditions ensuring the existence of a unique stationary solution to equation (1) are assumed (see, e.g., Chapter 3 in [2]). In particular, we consider that The marginals of are in the space of zero–mean –valued random variables with The autocovariance operator and the cross covariance operator are then nuclear operators.
The innovation process is assumed to be –valued Strong White Noise (–SWN) (see Definition 3.1 in [2]). Thus, has independent and identically distributed zero–mean –valued random components, with for all having functional variance which provides the trace norm of the autocovariance operator of
The following assumption establishes the AR(1) setting, where the results of the present paper are derived, and where, in particular, Carleman condition, and is satisfied by the respective marginals of and i.e., for and
Assumption A1. is a strictly stationary AR(1) process, having –SWN innovation such that with marginals and satisfying Carleman condition.
Under model (1), keeping in mind Assumption A1, the following orthogonality condition holds
for with denoting the almost surely (a.s.) equality.
2.1 Separating class in the Borel –algebra of
For a probability measure in its characteristic function is given by (see, e.g., Section 1.2.1 in [7])
Let be the marginal probability measure induced by and, for any Borel probability measure in consider
In what follows, denotes a centered non–degenerate Gaussian measure in with positive trace autocovariance operator
Lemma 1.
Under Assumption A1, if then,
Remark 1.
From Lemma 1, for any Borel measure in if for certain such that then with denoting the Borel –algebra of Then, one can conclude that any class of Borel sets of containing a set of positive –measure defines a separating class for the marginals of
In the introduction of a separating class of we consider the sets for a given orthonormal basis of with Note that
under Assumption A1. In the next lemma, the family of indicator functions will play a crucial role in the characterization of the orthogonality condition (2) for each where denotes the RKHS generated by the kernel of a positive trace operator Lemma 1(d) in [9] is now reformulated in our functional context.
Lemma 2.
Under Assumption A1, for each
where is given here in terms of the orthonormal basis of satisfying for a given positive trace operator on
Proof.
Let be the infinite product Gaussian measure on identified in sense with an infinite–dimensional Gaussian measure on with positive trace autocovariance operator satisfying (see Theorem 1.2.1 in [7]). Here, for each denotes a one–dimensional Gaussian measure with zero mean and variance For Hence, where denotes the standard normal cumulative distribution function. Thus,
with, as before, From Lemma 1, Remark 1 and equation (5), defines a separating class of Borel sets of This separating condition is sufficient for the class of functions to satisfy (4). ∎
For each the next lemma provides the martingale difference property of the –valued sequence
Lemma 3.
For each the sequence
is an –valued martingale difference with respect to the filtration where for
The proof is straightforward from for every with being the orthogonal projector into (see equations (2.55) and (3.4) in [2]).
2.2 Testing the autocorrelation model
For a given consider the testing problem:
Our formulation of K–S test is based on Lemma 2, and the –valued empirical process
Theorem 2 below establishes the weak convergence of to an –valued Gaussian process with covariance operator defined, for every as
Here,
From equations (7)–(8), process can be identified in law with time–changed –valued Wiener process with subordinator
That is, can be generated from –valued Wiener process on by inverse subordination via with
where now denotes the orthonormal basis of given by the eigenfunctions of the trace autocovariance operator Note that in our –valued context, infinite–dimensional tight Gaussian measures play the role of the uniform measure in finite–dimensions (see Theorem 4.1 in [6]).
Applying Theorem 2, and Continuous Mapping Theorem, we arrive at the following asymptotic approximation of K–S test in infinite dimensions:
as where denotes an –valued Wiener process on the interval with autocovariance operator in the sense introduced in Definition 2 of [8]. This process induces a probability measure on the separable Banach space, under the supremum norm of –valued continuous functions on with respect to norm. From (11), the critical values of the test based on can be approximated from the boundary crossing probabilities of
The implementation in practice of this testing criterion is complicated. Particularly, we are interested on arriving at an ADF test, which will be now introduced via random projection methodology. Then, we consider the test statistics
conditionally to the functional value of a Gaussian distributed functional random variable inducing measure Here,
Note that, from Theorem 2 below, under
conditionally converges in law to a Gaussian process on the interval with covariance function, given by, for
This process can be identified in law with real–valued time–changed Brownian motion with inverse subordinator Thus, as
where denotes real–valued Brownian motion on By Reflection Principle, the identity then allows the approximation of the critical values of this random projected based test statistics. As before, denotes the standard normal cumulative probability distribution function.
Under conditions of Lemmas 1–2, since, under Assumption A1, the marginals of also satisfy Carleman condition, the procedure can be summarized as follows: The null hypothesis holds if and only if, for with such that does not exceed a critical value of the probability distribution of the supremum norm of Brownian motion on Otherwise, exceeds a critical value of the probability distribution of the supremum norm of Brownian motion on –a.s. in which is equivalent to fails with probability one in view of Theorem 4.1 in [6].
Remark 2.
In practice, can be generated from an infinite–dimensional Gaussian measure Thus, can be generated from the identity
where and is a sequence of independent and identically distributed standard Gaussian random variables.
Note that equation (15) can be approximated in a consistent way from the results derived in Chapters 4 and 8 of [2]. Given the –SWN property of alternative strongly consistent approximations of the eigenvalues of follow from the strong law of large numbers. Specifically, for the case where is known, one can consider
While, in the case of being unknown, the empirical complete orthonormal eigenfunction system and the corresponding empirical eigenvalues of can be considered. They satisfy the following identity with for
3 Asymptotic probability distribution under simple
The main asymptotic results are derived in this section. Specifically, Theorem 1 in Section 3.1 provides, for each the convergence in distribution of the marginal to The random variable is distributed as a zero–mean –valued Gaussian random variable with autocovariance operator The weak convergence of to time–changed –valued Wiener process with subordinator introduced in equation (9), and covariance operator given in equation (7), is proved in Theorem 2 in Section 3.2.
Through this section we omit the dependence of the functional marks on the totally specified autocorrelation operator under
The next assumption is of technical nature, and plays a crucial role in the derivation of the results of this paper.
Assumption A2. Assume that is independent of for all
Remark 3.
3.1 Central Limit Theorem
Theorem 1.
Let be a centered AR(1) process, satisfying Assumptions A1–A2. Then, for each the following convergence holds:
with denoting the convergence in distribution, and, for each the autocovariance operator of the limiting Gaussian random variable is given by Thus, as
Proof.
The proof of this result consists of the verification of conditions (2.59)–(2.61) in Theorem 2.16 in [2]. Specifically, condition (2.59) follows from Vitali Convergence Theorem by proving convergence in probability to zero, and uniform integrability of Condition (2.61) provides tightness, and condition (2.60) requires the application of (2.36) in Corollary 2.3 in [2].
Let us first consider condition (2.59) in Theorem 2.16. For each we have to prove
Applying Chebyshev inequality, –SWN property of and strictly stationarity of under Assumption A1,
The Dominated Convergence Theorem yields convergence to zero of leading, from equation (19), to
Under Assumption A1,
ensuring uniform integrability. Thus, from equations (20) and (21), equation (18) follows from Vitali Convergence Theorem.
Let now prove, for each
corresponding to condition (2.61) in Theorem 2.16 in [2], where Specifically, for any
in view of the trace property of
Finally, we prove
corresponding to condition (2.60) in Theorem 2.16 in [2]. It is sufficient to prove condition (2.36) in Corollary 2.3 in [2], which is formulated in terms of Specifically, under strictly stationarity of and –SWN property of we obtain
where under Assumption A1, and under the trace property of From Corollary 2.3 in [2], with for all
for each and equation (23) holds.
3.2 Functional Central Limit Theorem
This section provides the convergence in law of to time–changed –valued Wiener process by applying an invariance principle based on Robbins-Monro procedure (see Theorem 2 in [29], which has been formulated in Lemma 4 in A.2.
Theorem 2.
Under Assumptions A1–A2, as the empirical process weak converges in distribution, in the space to a generalized –valued Gaussian process with covariance operator
for a given orthonormal basis of
Proof.
The proof of this result consists of the verification that the time–changed sequence of random elements in constructed in equation (26) below from –valued empirical process satisfies conditions (i)–(iii) of Lemma 4 (see Appendix). Specifically, equations (27), (28) and (29) below lead to the desired result, since the unitary jumps of go to zero in probability from (18), and tightness holds from (22). A Cramér–Wold device argument is applied to prove that all the finite–dimensional distributions of convergence to a multivariate infinite–dimensional Gaussian probability distribution with covariance matrix operator having functional entries given by equation (42) (see Appendix). Thus, the weak convergence in the space of to zero–mean time–changed –valued Wiener process with covariance operator
holds, with and for any
Let be the time–changed sequence of random elements in defined by
for each where time–changed is introduced by inverse subordinator and hence, (see equations (8)–(10)). As before, denotes the empirical process in (6), and is the –valued martingale difference sequence in Lemma 3.
From equation (40) in the Appendix, applying strictly stationarity and –SWN property of for any and
In particular, Lemma 4(ii) holds for
Finally, we prove uniform convergence in Lemma 4(iii).Specifically, for any
Since the Dominated Convergence Theorem leads to
where absolutely integrability criterion has been applied. Hence,
Equation (29) then converges to zero as uniformly in In particular, Lemma 4(iii) holds for
Equations (27), (28) and (29) also imply, in particular, the weak convergence, in the Skorokhod space of to an –valued zero–mean Wiener process with autocovariance operator of
3.3 Consistency
In the derivation of consistency of GoF test, the following assumption is considered, which is of technical nature, providing, in particular, information about the allowed distance between the null and the alternative hypothesis in the Hilbert–Schmidt operator norm.
Assumption A3. where denotes the autocorrelation operator under i.e., with being the autocorrelation operator under which is assumed to be known. For each we also assume that
In the next result, we will denote for
and we will consider for
Theorem 3.
Under Assumptions A1–A3, as
where means that expectation is computed under the alternative, and
with for
Proof.
The proof of this result provides an extension to the –valued context of the methodological approach adopted in pp. 217–218 of [23]. It follows from condition (2.36) in Corollary 2.3 in [2].
Under Assumptions A1–A3, to apply Corollary 2.3 in [2] we consider
Applying strictly stationarity of and equation (17), we obtain
for every where denotes the sequence of singular values of operator Its corresponding right and left eigenfunction systems are and Here, for
and
denotes the autocovariance operator of the truncated random variable is the power of operator which, as usual, it is understood as th self–composition of with since for invertibility of model (1) (see also Lemma 3.1, p.74 in [2] for more general formulations of such a condition).
Note that
and under Assumption A3,
Remark 4.
Conditionally to consistency of the test based on follows, in a similar way to (33), considering, for
and
4 Asymptotic properties for composite null hypothesis
In this section, we work under the assumption that is unknown, and must be estimated in a consistent way. We adopt the projection estimation approach introduced in Chapter 8 in [2], where two scenarios are distinguished, respectively corresponding to the cases of known and unknown eigenfunctions of the autocovariance operator of Here, we refer to the case of unknown eigenfunctions of operator of For the case of known eigenfunctions, see A.3.
Let be an AR(1) process satisfying the following conditions:
The eigenvalues of are such that
where, as before, for every
For every and consider and such that
and assume that a.s.
We consider the estimator given by
with and and for and
Assume that the conditions of Theorems 8.7–8.8 in [2] hold under the null
Hence, provides a strongly consistent estimator of under the null.
We consider the plug–in –valued empirical process
The following result derives the asymptotic equivalence in probability of and
Theorem 4.
Under conditions of Theorems 3.9 and 8.7–8.8 in [2], the following identity holds:
Proof.
The main ingredients in the proof of this result are the strong consistency of derived in Theorems 8.7–8.8 of [2], and the convergence to zero in probability of obtained from Theorem 3.9 of [2]. Both results are applied to obtain the convergence to zero in probability of in –norm, leading to their asymptotic equivalence in probability.
Applying Cauchy–Schwartz inequality we obtain for every
From Theorems 8.7–8.8 in [2], as converges in probability to zero. From equation (3.35) in Theorem 3.9 of [2], for
where and depend on and Hence, converges to zero in probability when Therefore, equation (38) converges to zero as Thus, and are asymptotically equivalent in probability in the norm of ∎
5 Final comments
In the context of functional residual marked empirical processes indexed by an –valued covariate, GoF asymptotics based on their random projection would require the application of a bivariate version of Theorem 4.1 in [6]. Our paper exploits inverse subordination for dimension reduction, and technical simplification purposes. Furthermore, our AR(1) setting involves mild conditions, allowing the application of Central Limit Theorems, and Functional Central Limit Theorems for –valued martingale differences sequences. This fact constitutes an important advantage of our approach leading to GoF asymptotic analysis in infinite–dimensions. The novelty in our equivalent formulation of the null hypothesis relies on our characterization of a Borel separating class in via infinite–dimensional non–degenerate Gaussian measures with trace autocovariance operator, whose RKHS defines the index set of the separating class. The identification of these measures with infinite product tight Gaussian cylindrical measures on allows us to work in a friendly environment, closely connected with finite–dimensional projections and cylindrical sets. Indeed, our approach is supported by the similar role that these Gaussian measures in infinite–dimensions play in relation to the uniform measure in finite dimensions.
In a near future, we also incorporate to our GoF asymptotic analysis, the formulation of ADF bivariate random projected tests, extending currently applied GoF asymptotic methodology in [5], for the functional linear model with scalar response, to the context of the functional linear regression model with –valued response and covariate. We will cover the cases of independent, and weak–dependent functional data. Another challenging topic to be addressed is the formulation, and asymptotic analysis of GoF tests for stationary LRD functional time series in the spectral domain, where the approach presented in [28] can support some advances, in the framework of functional time series in connected and compact two–point homogeneous spaces (see also [24] for the nonstationary case).
Acknowledgements
This work has been supported in part by projects PID2022–142900NB-I00, PID2024-158017NB-I00, and PID2020-116587GB-I00, financed byMCIU/AEI/10.13039/501100011033 and by FEDER UE, and CEX2020-001105-M MCIN/AEI/ 10.13039/501100011033), as well as supported by grant supported by grants ED431C 2021/24 and ED431C 2025/03 (Grupos competitivos) financed by Xunta de Galicia through European Regional Development Funds (ERDF).
Appendix
Appendix A Auxiliary information
Additional supporting information to facilitate the reading of the paper is provided in this section. Specifically, the second–order properties of –valued martingale difference sequence are provided in A.1. An invariance principle based on Robbins–Monro procedure is formulated in Lemma 4 in A.2. The limit infinite–dimensional Gaussian marginal distribution, and the limiting process of a plug–in functional residual marked empirical process indexed by , under known eigenfunctions of the autocovariance operator is analyzed in A.3.
A.1 Second–order properties of
Note that, for each the strongly integrability of the marginals of the –valued martingale difference sequence allows the computation of the means and conditional means of the elements of this sequence from their weak counterparts (Section 1.3 in [2]).
Under Assumptions A1–A2, for any and and can be computed as follows:
where we have applied for all under Assumptions A1–A2. As before, denotes the infinite–dimensional marginal probability measure induced by and is the event for each Keeping in mind equation (17), applying the strictly stationarity of for each the autocovariance operator of is given by
In a similar way, the covariance operator can be computed for every
where, for if and if
A.2 Invariance principle based on Robbins–Monro procedure
Lemma 4.
Let be a martingale difference sequence of –valued random variables, with respect to the filtration satisfying Let be a symmetric, positive semidefinite, trace operator. For each denote as the autocovariance operator of given That is,
Assume that
For
Then, the sequence of random elements in with the supremum norm, which are defined by
converges in distribution to a Brownian motion in with a.s., and covariance operator of
A.3 Composite null hypothesis under known eigenfunctions of
Let be an AR(1) process satisfying equation (1), and the following conditions:
is a standard AR(1) process with –SWN innovations, and
The eigenvalues of are strictly positive.
for every
Under the conditions assumed in Lemma 8.1, and Theorems 8.5–8.6 of [2], a consistent estimator of is given by
where and and
with being, as before, the eigenfunction system of which is assumed to be known.
Note that under (i)–(iii),
The design of K–S test under this scenario is based on the generalized –valued plug–in empirical process
Specifically, we consider the composite null hypothesis
where the family of AR(1) models under the null satisfies the conditions assumed in Lemma 8.1, and Theorems 8.5–8.6 of [2].
Appendix B Illustration of the performance of GoF
This section presents two illustrations of the proposed GoF test in the context of AR(1) (), and SPAR(1) models in B.1 and B.2, respectively. In the last case (see, e.g., [3]), the assumed invariance property of the kernels defining the autoregression, and covariance operators of SPAR(1) process leads to an important dimension reduction in the implementation. Similar results hold under this invariance property in the case of autoregressive processes on compact and connected two point homogeneous spaces (see, e.g., [25]; [26]; [27]).
B.1 Detecting independence and nonlinearities
Finite sample performance of the proposed GoF test is now illustrated under simple and composite hypothesis.
Let us first consider the simple null hypothesis
For comparative purposes, in this section, a bit different (from (12)) random projection implementation is performed, adopting the numerical approach in [1] and [11]. The type I empirical error, and the empirical power, based on repetitions, for functional samples sizes are computed. The critical values of the test statistics are approximated from Fast Bootstrap based on bootstrap replicates.
Under a Gaussian scenario, the functional values of an AR(1) process are generated with support in the interval evaluated at temporal nodes. The innovation process is assumed to be Gaussian –SWN noise. Its integral autocovariance operator has exponential kernel given by
A Gaussian random initial condition is also generated, independently of the innovation process . has exponential autocovariance kernel having the same scale parameter values, as the innovation process . The recursive equation (1) is implemented from the numerical approximation of the integral where, under the alternative , the kernel of the integral autocorrelation operator is given by for . The constants are ensuring that the process is stationary. To remove dependence from the random initial condition, a burn-in period of observations is considered. That is, for each functional sample size the above–described recursive approximation of the values is computed removing the values
As indicated, for comparative purposes, the random projection methodology is implemented from a bivariate infinite–dimensional Gaussian measure with independent components. Thus, we consider the bivariate random projected empirical process
where and are independent realizations of zero–mean Gaussian processes with autocovariance operators and respectively. These processes are generated from their truncated, at term Karhunen–Loéve expansions.
Fast Bootstrap is implemented from the equation
where are independent and identically distributed standard normal random variables. The computation of the –value is obtained for each projection from
where, as commented, bootstrap replicates have been considered. The numbers of projections tested is To obtain only one –value the False Discovery Rate is computed (i.e., the expected proportion of false positives among the rejected hypotheses).
For the case of composite null hypothesis the process generated is an AR(1) process, that uses the same employed to the alternative hypothesis in the case of simple null hypothesis. For the alternative, we have also considered an AR(1) process with the same but evaluated in the square process, i.e.
where the parameter avoids that the process escapes to nonstationarity, being only when . Although, this is a clearly non AR(1) process, it is pretty close to an AR(1) due to the scale of . Now, the statistics in equation (47) must change to incorporate the variability due to the estimation of parameter . So, following [9], is changed by given by
where is the estimation of the functional parameter and with being the resampled errors from the estimated model.
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.050 | 0.054 | 0.044 | 0.038 | 0.034 | 0.034 | 0.042 | |
| 0.060 | 0.056 | 0.054 | 0.062 | 0.058 | 0.060 | 0.056 | |
| 0.050 | 0.054 | 0.056 | 0.052 | 0.048 | 0.052 | 0.040 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.434 | 0.492 | 0.526 | 0.524 | 0.530 | 0.562 | 0.552 | |
| 0.486 | 0.630 | 0.678 | 0.728 | 0.764 | 0.836 | 0.848 | |
| 0.650 | 0.808 | 0.896 | 0.922 | 0.932 | 0.984 | 0.994 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.054 | 0.034 | 0.034 | 0.048 | 0.056 | 0.038 | 0.032 | |
| 0.056 | 0.062 | 0.058 | 0.050 | 0.068 | 0.046 | 0.046 | |
| 0.050 | 0.050 | 0.052 | 0.046 | 0.054 | 0.048 | 0.048 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.114 | 0.096 | 0.084 | 0.062 | 0.064 | 0.062 | 0.066 | |
| 0.152 | 0.182 | 0.196 | 0.204 | 0.214 | 0.232 | 0.264 | |
| 0.242 | 0.340 | 0.378 | 0.400 | 0.432 | 0.520 | 0.508 | |
| 0.302 | 0.438 | 0.528 | 0.562 | 0.584 | 0.666 | 0.706 | |
| 0.448 | 0.610 | 0.680 | 0.756 | 0.806 | 0.892 | 0.938 | |
| 0.474 | 0.682 | 0.774 | 0.830 | 0.888 | 0.970 | 0.992 |
In Tables 1 and 2, one can respectively find the –values and power approximations computed, based on repetitions of the implemented testing procedure, for functional sample sizes The Confidence Interval, based on repetitions, obtained for is Specifically, Table 1 shows the rejection probability when is true. One can observe in Table 1 that the computed estimates of the test size are slightly below the true parameter value for the functional sample size . While, for functional sample sizes an improvement is observed in the computed estimates of for all number of random projections. For the functional sample sizes analyzed, no patterns are observed regarding computed estimates and the number of random projections. The empirical power of the test, for the same functional sample sizes and number of repetitions, are displayed in Table 2. As expected, the performance of the test is improved when the functional sample size increases. Tables 3 and 4 show sizes and powers for the composite hypothesis. The test seems well calibrated as every element in Table 3 is inside the Confidence Interval for although, for the test shows a tendency to be below similar to what happens in simple null hypothesis. Given the rate of convergence of the empirical eigenfunctions to the theoretical ones (see Lemma 4.3 and Theorem 4.5 in [2]), larger functional sample sizes are required to obtain competitive empirical power values. In Table 4, we illustrate this fact considering additional functional sample size values
B.2 Spherical functional time series SPAR(1)
In this section we analyze the performance of the proposed GoF procedure in the special case where is the space of real–valued square integrable functions on the –dimensional sphere in Here, denotes the normalized Riemannian measure on In what follows, we will denote by the orthonormal basis of eigenfunctions of the Laplace–Beltrami operator on with denoting the dimension of the th eigenspace of the Laplace–Beltrami operator. Here, we consider
The time–varying random projections of the SPAR(1) process, with repect to the eigenfunctions of the Laplace Beltrami operator, have been generated by using the MatLab function arima. They are evaluated at time instants, where denotes the functional sample size. The functional values of the innovation process are also generated from its time–varying projections with respect to such a basis at times.
In the generations of SPAR(1) model we have considered the spherical harmonics which are localized in the first four eigenspaces () of the Laplace Beltrami operator on (see left–hand–side of Figure 1). Under the simple null hypothesis one realization (evaluated at spherical spatial nodes) of the SPAR(1) process projected into the corresponding direct sum of Laplace Beltrami eigenspaces is displayed at the right–hand–side of Figure 1. Here, denotes the th eigenspace of the Laplace Beltrami operator on The eigenvalues of the invariant kernel defining the autoregressive operator under simple associated with the Laplace Beltrami eigenfunctions, are given by
We first consider versus where we assume that the response is of the form with for and being an –SWN. Thus, the random components of the innovation process still display significative correlations induced by Both operators, and are invariant bounded linear operators against the group of rotations in the sphere. The test statistics (12) is evaluated at different random projections generated from –valued Fractional Brownian Motion (FBM) with autocovariance operator The critical value corresponding to is for the supremum norm of Brownian motion on the interval As indicated, this value is computed, by applying Reflection Principle, from the standard normal probability distribution (see Section 2.2).
Under simple null hypothesis (totally specified autocorrelation operator), Table 5 displays the empirical test size for the functional sample sizes Again, no patterns are displayed. In particular, no tendency to be below is observed for In Table 6, one can find the empirical test power values, based on repetitions, for the functional sample sizes As in the previous scenario analyzed in Table 2, one can observe in Table 6 a better performance of the test statistics when the functional sample size increases.
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.060 | 0.040 | 0.040 | 0.060 | 0.040 | 0.040 | 0.060 | |
| 0.040 | 0.040 | 0.040 | 0.060 | 0.060 | 0.060 | 0.040 | |
| 0.050 | 0.054 | 0.056 | 0.052 | 0.048 | 0.052 | 0.040 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.376 | 0.336 | 0.718 | 0.454 | 0.542 | 0.594 | 0.584 | |
| 0.566 | 0.464 | 0.658 | 0.700 | 0.428 | 0.450 | 0.762 | |
| 0.562 | 0.772 | 0.806 | 0.772 | 0.744 | 0.882 | 0.732 | |
| 0.838 | 0.546 | 0.682 | 0.944 | 0.864 | 0.912 | 0.892 | |
| 0.890 | 0.998 | 0.914 | 0.994 | 0.978 | 0.984 | 0.984 | |
| 0.851 | 0.926 | 0.974 | 0.938 | 0.850 | 0.996 | 0.992 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.018 | 0.014 | 0.014 | 0.022 | 0.028 | 0.022 | 0.022 | |
| 0.032 | 0.034 | 0.026 | 0.022 | 0.034 | 0.030 | 0.024 | |
| 0.038 | 0.052 | 0.032 | 0.042 | 0.040 | 0.038 | 0.040 | |
| 0.042 | 0.054 | 0.032 | 0.026 | 0.056 | 0.040 | 0.042 | |
| 0.046 | 0.046 | 0.050 | 0.054 | 0.046 | 0.044 | 0.044 | |
| 0.050 | 0.045 | 0.050 | 0.045 | 0.045 | 0.045 | 0.055 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.636 | 0.706 | 0.522 | 0.506 | 0.484 | 0.584 | 0.576 | |
| 0.668 | 0.736 | 0.630 | 0.652 | 0.556 | 0.806 | 0.678 | |
| 0.824 | 0.870 | 0.646 | 0.912 | 0.612 | 0.062 | 0.638 | |
| 0.962 | 0.826 | 0.670 | 0.902 | 0.882 | 0.914 | 0.764 | |
| 0.994 | 0.966 | 0.998 | 0.978 | 0.986 | 0.918 | 0.840 | |
| 0.998 | 0.986 | 0.996 | 0.826 | 0.786 | 0.976 | 0.960 |
Let us now consider, under the general scenario given by condition (c0) in [2], ensuring existence and uniqueness of a stationary solution to the AR(1) equation (1) (see [2], p.74, Chapter 3), the following testing problem:
where generations under have been achieved for with eigenvalues Hence, under Generations under have been performed, as before, under the pure point spectrum ensuring is satisfied, since Note that Lemma 8.1(3) in [2] holds under the generated scenario, since for and computed from terms of the empirical sequence Here, denotes the nearest positive integer.
Tables 7 and 8 respectively show the empirical test sizes and powers for composite null hypothesis under a SPAR(1) scenario. For the number of random projections, functional sample sizes and number of repetitions reflected in Tables 7 and 8, the values of the test statistics (12) based on the plug–in empirical process (46) are respectively computed under and under The eigenvalues of are estimated from the following identity (see equation (3.13) in [2]):
where denotes the adjoint of with in the generations. Given the diagonal spectral factorization in terms of the common resolution of the identity of operators in equation (49), the associated pure point spectra satisfy
Equation (50) is then approximated in terms of the following empirical pure point spectra:
where, for and denote the generated repetitions of and under respectively, for the estimation of under in the SPAR(1) model. Note that, from equations (50)–(51), considering in the a.s. convergence in –norm in Lemma 8.1(3) in [2], the a.s. convergence of the eigenvalues of (computed from (50)–(51)) to the eigenvalues of satisfying (50) is obtained.
In Table 7, we observe empirical test sizes close to the theoretical value for the sample sizes The empirical test powers showed in Table 8 are also computed for the functional sample sizes considering the same number of repetitions As expected, the empirical test powers are improved when the functional sample size increases. Under an SPAR(1) scenario, a better performance in the case of composite null hypothesis is observed for small sample sizes since we work under a lower level of misspecification of the autocorrelation operator than in Table 4 for Under simple the reverse situation is observed in Tables 2 and 6. Although, given the variability displayed by columns in these two tables, for the sample sizes one can not conclude from these numerical results any significative performance difference between these two testing procedures.